40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 582     Entry time: Fri Jun 27 14:36:39 2008
Author: John 
Type: Summary 
Category: Locking 
Subject:  
Rob, Yoichi, John

Some progress last night:

Switched back to SPOB_DC for CARM.

Shaped the MC LSC loop to reduce excess noise in the 20-30Hz band. Likely the most significant change.
Could this be due to fan noise from the laptop on the SPOB table?

Brought in the AO path earlier (at low gain).

Reduced offset to 6 and increased MC gain before handing off to SPOB. Ramped up AO and MC gain then switched off the
moving zero.

Looks like PD11 is the most promising candidate for RF CARM although the demod phase needs attention.
Attachment 1: gettingcloser.png  108 kB  Uploaded Fri Jun 27 15:39:03 2008  | Hide | Hide all
gettingcloser.png
ELOG V3.1.3-